OpenStudy (anonymous):

If f(x) + x^2 * f(x) = 10 and f(1)=2, find f'(1).

OpenStudy (anonymous):

check ur ques. again its wrong it does not satisfy for f(1)=2 in the main equation

OpenStudy (anonymous):

well that's the friggen problem verbatim from the textbook

OpenStudy (anonymous):

OH pellet

OpenStudy (anonymous):

it's not sorry geez sorry it's f(x) + x^2 * f(x)^3 = 10 the second f(x) term is cubed

OpenStudy (anonymous):

Maybe rearrange the problem a bit....\[y+x^2y=10\implies y=\frac{10}{1+x^2}\]I don't think your problem meets the specified conditions.

OpenStudy (anonymous):

is ur anwer 2/5??

OpenStudy (anonymous):

it's -16/13 that's the answer in the book

OpenStudy (anonymous):

animalain - it's y + (x^2)(y^3) = 10

OpenStudy (anonymous):

Yeah, I got that. Trying to figure out my strategy....

OpenStudy (anonymous):

yeah its done

OpenStudy (anonymous):

...

OpenStudy (anonymous):

How did you do it?

OpenStudy (anonymous):

differentiate it n then put x=1 in equation u'll get the result

OpenStudy (anonymous):

is that done??

OpenStudy (anonymous):

\[y + (x^2)(y^3) = 10\implies y'+2xy^3+3x^2y^2y'=0\]\[\implies y'(1+3x^2y^2)=-2xy^3\implies y'=\frac{-2xy^3}{1+3x^2y^2}\]\[\implies y'(1)=\frac{-16}{13}\]Got it. Need to practice my calculus a little more....LOL

OpenStudy (anonymous):

all the best..:)